OpenStudy (anonymous):

Is this right? I am trying to calculate concentration. 10mm (liquid) to 7mm water = 70% @.Sam.

OpenStudy (anonymous):

thank you zach

OpenStudy (anonymous):

70% of what?

OpenStudy (anonymous):

of water

OpenStudy (anonymous):

Concentration of water in liquid (percent)

OpenStudy (anonymous):

ok well that would be true if adding them together made a total of 10mm of (liquid i guess) but adding them together gives you a total of 17mm not 10

OpenStudy (anonymous):

so yuo r fraction should be \[\frac{ 7 }{ 17 }\]

OpenStudy (anonymous):

I was given a chart with 10 mm amount of liquid and 7mm amount of water then to figure out the calculation in percentage of concentration of water in liquid. It's for science.

OpenStudy (anonymous):

I came out with 70%

OpenStudy (anonymous):

ok i might be wrong about the fraction

OpenStudy (anonymous):

so you are adding the water into the liquid no?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok and after we do that the jar/cup will have 17mm of liquid in it right?

OpenStudy (anonymous):

by liquid i mean the combination of water and the other liquid before we mixed them

OpenStudy (anonymous):

ok so it is 17% ? ?

OpenStudy (anonymous):

oh no ! Ooops I meant 7 not 17 way above sorry.

OpenStudy (anonymous):

numbers should be 10 and 7 not 17. Sorry : (

OpenStudy (anonymous):

well im sorry but i think they are both wrong. i set up a few fractions and got 41.2% water

OpenStudy (anonymous):

\[\frac{ 7 }{ 17 }=\frac{ x }{ 100 }\]

OpenStudy (anonymous):

cross multiply

OpenStudy (anonymous):

why 17?? Where did you get that? Did you add 10 + 7 you mean?

OpenStudy (anonymous):

yes let me put it this way...

OpenStudy (anonymous):

\[\frac{ water }{ water+liquid }=\frac{x }{ 100% }\]

OpenStudy (anonymous):

10mm total amout of liquid to 7mm water = 41.2 %

OpenStudy (anonymous):

ok writing it down.

OpenStudy (anonymous):

well the total amount of liquid is 17mm and if you picture 7+10 the 7 is going to be less than 50%

OpenStudy (anonymous):

I got it I understand ! Thank you so much !!

OpenStudy (anonymous):

no problem good luck with science!

OpenStudy (anonymous):

I'm calculating the others now. Thanks!